OpenStudy (anonymous):

create your own quadratic equation that has an axis of symmetry of x = 2 and whose graph open down. Then find the vertex, domain, and range and x-intercepts. show all work to receive full credit @uri @Opcode @ganeshie8 @.Sam.

OpenStudy (anonymous):

I got -2x^2 + 8x +5 is that right @uri @Opcode @ganeshie8 @.Sam.

OpenStudy (campbell_st):

well that opens down and has an axis of symmetry at x = 2 to check use x = -b/2a or x = -8/(2*-2) so to find the vertex, subsitute the line of symmetry value, x = 2, into the equation to find y